### v1

Abstract
The problem of using structured methods to represent knowledge is well-known in conceptual modeling and has been studied for many years. It has been proven that adopting modeling patterns represents an effective structural method. Patterns are, indeed, generalizable recurrent structures that can be exploited as solutions to design problems. They aid in understanding and improving the process of creating models. The undeniable value of using patterns in conceptual modeling was demonstrated in several experimental studies. However, discovering patterns in conceptual models is widely recognized as a highly complex task and a systematic solution to pattern identification is currently lacking. In this paper, we propose a general approach to the problem of discovering frequent structures, as they occur in conceptual modeling languages. As proof of concept for our scientific contribution, we provide an implementation of the approach, by focusing on UML class diagrams, in particular OntoUML models. This implementation comprises an exploratory tool, which, through the combination of frequent subgraph mining algorithm and graph manipulation techniques, can process multiple conceptual models and discover recurrent structures according to multiple criteria. The primary objective is to offer a support facility for language engineers. This can be employed to leverage both good and bad modeling practices, to evolve and maintain the conceptual modeling language, and to promote the reuse of encoded experience in designing better models with the given language.
keywords:
Conceptual Modeling, Mining Conceptual Models, Frequent Subgraph Mining, Recurrent Modeling Structures, Modeling Patterns1 Introduction
Conceptual modeling is a highly complex task. As empirical results show [1, 2, 3, 4, 5], this is often due to multiple factors, such as the cognitive limitation of modelers, the lack of information about the domain to be modeled, and the nature of the conceptual modeling language being adopted. For this reason, one of the main concerns of language engineers is overcoming modelers’ difficulties and devising conceptual modeling languages that assure as much as possible the high-level quality of the output conceptual model. Patterns play a key role in this regard. These recurrent structures are generalizable solutions to design problems that help in understanding and improving the process of creating models. Specifically, patterns represent examples of good (or bad) recurring practices that language engineers can identify by looking at the practical application of the language itself. By discovering and adopting recurrent structures, language engineers can then evolve the modeling language itself, speeding up and facilitating its application and easing the reuse of working experiences, by also helping to avoid common errors or misconceptions.
During the past decade, modeling patterns have been widely adopted by language designers [6, 7] and, consequently, discovering them from design experiences has become of paramount importance. Still, the task of discovering recurring modeling structures across conceptual models presents multiple challenges. For instance, consider that .i recurrent modeling practices can only be discovered by observing a vast number of conceptual modeling examples, and .ii the assessment of the conceptual models and the identification of recurrent modeling structures often require analysis activities that are highly time-consuming when performed manually, as seen in processes like modularization, frequency calculation, or constructs correlation.
Recently, approaches to facilitate this complex discovery process produced valuable results [8, 9]. An empirical study [10] demonstrated that the Frequent Subgraph Mining algorithm [11] can be used to automatically discover certain kinds of patterns in OWL vocabularies. Works like [8] showed how the implementation of Tree Mining [12] techniques aids the detection of specific types of patterns in BPMN. Still, a general solution to this problem has not been devised.
In this paper, we present an approach for facilitating the empirical discovery of recurrent modeling structures. We implemented this approach in a tool that acts as an exploratory facility, which allows users to discover and analyze patterns interactively. The tool relies on the application of gSpan [13], a well-known frequent subgraph mining algorithm. Through a command-line interface, it allows users to:
.i
select the models to be processed,
.ii
prepare the data to be mined by focusing on specific information and filtering out concepts that might be not relevant for that type of analysis,
.iii
select certain features that the output patterns should have (e.g., number of nodes, frequency, or dissimilarities from known patterns),
.iv
visualize the output patterns and query the input models to verify the domain occurrences of that pattern, and
.v
cluster the output patterns to simplify the user’s final assessment.
The main scope of our solution is to offer a support facility for language engineers that aims at exploiting bad/good practices to evolve and maintain the conceptual modeling language, but also to favor the reuse of encoded experience in designing better models with the given language.
We evaluate our approach using a dataset of UML Class Diagrams, in particular OntoUML models. We adopted this language for several reasons. Primarily, OntoUML is a language based on pre-defined patterns, which are frequently used across different models to address common modeling issues. This can allow us to evaluate our approach to finding useful structures that are already known by the designers of the language itself. Second, for our evaluations, we can access a curated dataset of OntoUML models [14], encoded in a uniform format. Third, for our evaluations, we can obtain feedback directly from the authors of the language, enabling us to gather valuable information about the practical utility of the approach. Finally, OntoUML has a variety of constructs and features that allow us to assume that if our approach works for this language, it also works for widely used standard languages, such as UML or BPMN.
The structure of this paper is as follows: Section 2 explores the concept of “pattern” in conceptual modeling, focusing particularly on “recurrent modeling structures”. This section also provides a brief overview of the primary technique employed for automating the mining process. Section 3 enumerates the requirements that guided the design of our methodology. In Section 4, we introduce our framework, elaborating on its workflow and individual components. Section 5 details the implementation of our approach. In Section 6 and Section 7, we discuss the experiments and demonstrations conducted to validate our solution. In Section 8, we analyze insights gained from the evaluation and demonstration processes. Section 9 situates our work within the context of existing literature. Finally, Section 10 offers reflections on our findings.
2 Research Baseline
2.1 Patterns in Conceptual Modeling
Currently, our proposal primarily focuses on patterns in conceptual models. These model fragments represent recurrent structures formed by modeling constructs from a specific conceptual modeling language. Keeping track of these patterns allows one to understand how a language is applied, thus offering powerful means to language designers for maintaining and evolving the language itself. Using recurrent structures, one can, e.g., discover the most common combinations of language constructs, identify language dialects for specific application domains, verify possible restrictions in using constructs or their combinations, and determine the frequent subversions of the language.
The notion of systematic language subversion [15] refers to an ungrammatical use of a language’s constructs that becomes recurrent in a language community signalling a design limitation of that language. It is closely related to coding traditions [16] that cover coding policies, notational guidelines, naming conventions, implementation patterns, programming idioms, etc.
Figure 1 shows examples of common usages of the generalization relationship in UML class diagrams. These structures represent the ways by which modelers represent intentional sub-typing (inheritance) between classes (see .a, .b and .c types), and can be taken as useful insights for understanding the language application. For instance, given a set of models about a certain domain of information, it is possible to discover that the multiple inheritance (.c) is used to capture class compositionally [17]. This observation might trigger language designers to suggest class composition relations when modelers use two or more generalization relations for a class.
Similarly, Figure 2 represents three examples of recurring structures in BPMN models. These patterns correspond to equivalent ways of modeling tasks parallel splitting (see .a, .b, and .c types) and are related to how people design control flows. Just as with UML structural patterns, these BPMN fragments can offer interesting insights into the language’s usage. For instance, by observing multiple occurrences of the implicit parallel split (.b), language designers may decide to force modelers to make explicit the gateway for the splitting, to avoid unintended instantiations of the model (i.e., where the splitting may be intended as an ‘‘OR’’, instead of an ‘‘AND’’, depending on the modelers’ scope).
2.2 Frequent Subgraph Mining
Frequent Subgraph Mining (FSM) is a well-known technique [11] used to find frequent subgraphs in a graph dataset. This technique typically involves a dual-phase approach. The initial stage involves the creation of subgraph candidates [11], where subgraphs in the graph dataset are searched and proposed for the analysis, while the subsequent step entails evaluating the frequency of the generated subgraphs to ascertain their prevalence.
FSM presents two distinct problem formulations: .i graph transaction-based FSM, and .ii single graph-based FSM. In graph transaction-based FSM, the input dataset comprises a collection of medium-sized graphs termed “transactions” [19]. On the other hand, single graph-based FSM uses a single, significantly large graph as input. In this setting, a subgraph, denoted as g, is deemed frequent if its occurrence count surpasses a predefined threshold value, commonly referred to as the support threshold. The support calculation varies according to the problem formulation. In a transactional dataset, the support of a graph is the ratio of the number of transactions to which this graph occurs to the total number of transactions. In a single large graph, the support of a graph is the number of its occurrences in this dataset.
Formally, given a graph , where is a set of (possibly labeled) vertices and is a set of (possibly labeled) edges, then a graph is a sub-graph of G if and only if its vertices and edges are a subset of the vertices and edges of graph G. A subgraph is considered frequent when its support equals or exceeds the user-defined minimum support threshold.
Let us consider a graph database represented in Figure 3. Suppose that we want to discover all the (connected) subgraphs with at least two nodes that occur in at least three graphs. By adopting minimum support 3 as a parameter and applying an FSM algorithm, we can obtain the set of all subgraphs appearing in at least three graphs, namely .
In practical scenarios, different tools or algorithms can identify frequent subgraphs. In this paper, we adopt gSpan [13], a popular and important state-of-the-art solution. This algorithm is often used for discovering frequent subgraphs in large graph databases, and has variants that work on conceptual graphs [20], typed linked data [21], chemical compound data [22], large disk-based databases [23], data with differential privacy [24], etc. It uses an approach based on a minimum DFS code [25] that allows to efficiently generate candidate subgraphs and prune infrequent ones. The algorithm iteratively discovers all frequent subgraphs containing an edge and then expands the search to include larger subgraphs. The database is shrunk as the search continues, and only graphs containing the current subgraph are considered. Studies show that gSpan can mine large frequent subgraphs with lower minimum supports and high performance [11].
3 Requirements
Our design for the proposed approach is based on an initial problem identification activity. During this phase, we gathered feedback from five potential users of our approach: expert language creators, who have been involved in the development of different conceptual modeling languages and who have been working on the identification of conceptual modeling patterns. We performed open-ended interviews and the main open questions we asked were:
.i
What is the relevance of an approach for facilitating the empirical discovery of structural modeling patterns in conceptual models?, and
.ii
What is required in order to facilitate the empirical discovery of modeling patterns?
This preliminary step helped us improve our awareness of the problem, better understanding the related work, and better identifying the features that our solution should offer to the end-users. The feedback from experts was crucial in defining both functional and non-functional requirements, which are needed to design the approach and evaluate the artifact in which our contribution is embedded.
We mapped the key features that specify what our approach should do into the following functional requirements:
R1.
Interestingness: The approach should facilitate the discovery of subjectively interesting patterns, namely recurrent structures that can be considered interesting according to the user’s interpretation. Here, the notion of “subjectively interesting” is inspired by the work from Silberschatz and Tuzhilin [26], where a pattern is ranked as interesting by a user mainly because .i it is considered exploitable for the modeling activities, or .ii it contradicts some user’s expectations.
R2.
Customization: The approach should allow engineers to manipulate the input models to obtain different patterns, which may vary in size (number of nodes or edges), may be related to different constructs (see, for instance, taxonomical vs. non-taxonomical structures), or maybe taken at different levels of granularity (e.g., in UML compositions and aggregations relations may be taken just as associations). The approach should also allow the user to filter the output of the discovery process according to custom parameters.
R3.
Comprehension: The approach should assist engineers in the process of assessing and analyzing the output patterns. This should be feasible by accounting for multiple types of frequency, e.g., relative frequency, as the number of occurrences of a pattern in each model; total frequency, as the overall occurrences for a given pattern; frequency across models, as the number of models in which a pattern occurs. Moreover, this should be supported by a human-readable visualization format of the output patterns.
We mapped the key features that specify how the approach should perform its functions into the following non-functional requirements:
R4.
Performance. The approach should outperform the human discovery activity in terms of time. Moreover, the conceptual models processing and the mining step should happen in a reasonable amount of time, even with a large amount of data, where “reasonable” and “large” are considered concerning related work [27, 28] (e.g., hundreds of models).
R5.
Compatibility. The approach should be able to support pattern discovery in multiple conceptual modeling languages.
4 Method
Our approach is represented as a workflow, consisting of several tasks, which can be categorized into three main phases: preparation, discovery, and assessment. These phases allow the user to intervene in the workflow whose inputs, outputs, and dependencies are combined as from Figure 4 below.
4.1 Preparing the Input Data
The “Preparation” phase focuses on transforming the input conceptual models in a format that is processable by the subsequent steps. In this phase, we have two main tasks.
First, the Importing (0) task consists of taking a set of conceptual models encoded in a language (e.g., UML or BPMN) and transforming each model into a Labeled Property Graph (LPG) .111Note that this task is language-dependent and requires an ad hoc transformation for each source conceptual modeling language. At first glance, this might seem straightforward since conceptual models are essentially graphs. That, however, is not always the case. Consider, for instance, the transformation of UML class diagrams into graphs. The simple solution is transforming classes into nodes and generalizations and associations into edges. Still, if we want to convert, generalization sets, association classes, generalizations between associations, cardinalities, and several other constructs, this solution no longer works. Moreover, a challenge for the available frequent subgraph mining approaches is accounting for complex (semantically-rich) graph data [11, 29, 30]. There are several algorithms each of which admits graphs with different characteristics (e.g., labeled vs. unlabeled or directed vs. undirected). In our case, we have to account for highly expressing modeling languages, such as OntoUML, which represent limiting cases that challenge the functionality of the available mining solutions. Therefore, we opted for full compatibility with all kinds of algorithms, solving the problem at the source and generating a format that is always processable.
Data: Set of of Conceptual Models Result: Set of Labeled Property Graphs (LPG) 1 for every conceptual model , where do 2 map each class into a node ; 3 map each association into a anode ; 4 map each generalization into a node ; 5 map association cardinalities into a node ; 6 for each generalization node do 7 Connect to the parent class-node with an edge “general”; 8 Connect to the child class-node with an edge “specific”; 9 end for 10 for each association node do 11 Connect to the source class-node with an edge “source”; 12 Connect to the target class-node with an edge “target”; 13 for each cardinalities node do 14 Connect to the corresponding with an edge “cardinalities”; 15 end for 16 17 end for 18 for each class do 19 create the corresponding label(s) for node ; 20 end for 21 for each association do 22 create the corresponding label(s) for node ; 23 24 end for 25 add a graph to ; 26 27 end for 28
Algorithm 1 offers a partial view of how the importing step works. The main rationale is then to reify everything that is in the input conceptual model. For instance, associations, cardinalities, and classes are all mapped into graph nodes with multiple labels. Figure 2 provides an example of transformation occurring via importing. .a represents the input conceptual model, encoded into OntoUML, .b represents the output labeled property graph. The edge labels “source-target” and “specific-general” are adopted to keep track of the directionality of generalizations and associations, respectively.
In the Filtering (1) task, users can choose which language constructs to exclude from the models. For instance, with OntoUML, one may want to seek patterns only involving classes decorated with certain stereotypes, or involving only classes, generalizations, and generalization sets. Finally, the outputs of the filtering task are a set of (possibly filtered) LPGs (.ii in Figure 4) to be given as input to the mining algorithm and the storage of all the generated graphs (.iii, which can be used to run the Deepening (3) task (see Section 4.2 below).
4.2 Discovering Patterns Across and Within Models
The “Discovery” phase is the core of the workflow and is aimed at generating the candidate patterns in a format that can be then processed and made accessible to the user for the final assessment. This phase, in turn, is composed of two main tasks.
Mining (2), the first task, involves applying an FSM algorithm. This allows for an interaction with the user, who can select:
.i
the minimum pattern size for the output patterns (e.g., filter out patterns that have less than five nodes),
.ii
the minimum frequency threshold for the output patterns (e.g., filter out patterns that occur less than 30 times across models), and
.iii
the known patterns to be excluded from the final output (e.g., the user may want to filter out subgraphs that are highly expected to focus on unexpected patterns).
The output of the mining task will then comprise a list of the discovered patterns in a format that eases the final assessment, along with a pattern index, a list of indexes of the source models in which the pattern occurs, and a model frequency value calculated as the number of conceptual models that contain the pattern. For instance, given five models, the model frequency cannot be more than ‘5’.
An optional task named Deepening (3) allows users to select a pattern from the previous task and discover the occurrences of this pattern within each input model. This operation can be run for each of the output patterns and implies using the selected pattern as a query to be run over the graphs storage generated via the importing task. The output of the deepening task is a set of pattern occurrences for each selected pattern. This allows the user to derive the total frequency for each pattern, calculated as the sum of the pattern occurrences within each input model (e.g., given five models, if the pattern occurs twice in each model the total frequency will be of ‘10’, but the model frequency will be ‘5’). Through this step, the user can extract information that is related to the specific usage of the pattern within the reference model domain. For instance, via deepening, we can infer that, in a pattern, a node with stereotype Kind is associated with the label ‘Vehicle’ in a model and ‘Means of transportation’ in another model.
4.3 Assessing the Output Patterns
The “Assessment” phase aids users in analyzing the output. Similar to previous phases, this one comprises two main tasks.
Data: Set of of Frequent Structures Result: Set of Clustered Frequent Structures 1 for every frequent structure , where do 2 extract as the number of nodes in ; 3 extract the number of edges in ; 4 extract the adjacency matrix for ; 5 for each node where do 6 extract the associated labels ; 7 end for 8 for , , and do 9 flatten into a vector concatenating its rows; 10 generate a vector ; 11 store in a list of vectors ; 12 end for 13 14 end for 15for every vector do 16 compute similarity with every other vector ; 17 take as input a similarity threshold ; 18 create the set of clusters according to 222Note that, in our scenario, if we have three patterns, A, B, and C, and, according to a certain threshold, A is similar to B and B is similar to C, this implies that A is also similar to C, and A, B and C are in the same cluster.; 19 20 end for
The Clustering (4) task groups output patterns—specifically, the recurrent or frequent structures from the mining task—based on their similarities.333Notice that in the approach we are proposing, to assess whether one conceptual model sub-structure is close to another, we take inspiration from conceptual model similarity techniques [31, 32], where the main task is to find similar models given a reference model input, to categorizing models according to their characteristics. This should facilitate the user’s process of scraping and consulting the output. Often, the number of outputs produced can surpass hundreds, and numerous patterns can have similar characteristics. For example, some structures might differ by just one label (in one structure we might have three nodes where one is labeled as Kind and the other two specialize it as Subkind, while in another structure we might have three nodes where one is Kind, and the other two specialize it as Subkind and Role, respectively. In this sense, a cluster of patterns can be considered by the user as a set of variations, which sometimes may refer also to the same structural pattern (e.g., where a kind is specialized by two or three Subkinds). To enable clustering, in our process, we extract key characteristics from the graphs that represent each pattern. At present, the steps addressed to enable the clustering step are as from Algorithm 2 below.
Currently, we adopt a single feature extraction method to calculate pattern similarity. However, this does not prevent adopting different feature extraction or embedding techniques. This is valid also for the approach used to assess similarity, which now is limited to a cosine similarity [33], but in the future, it can be extended to other measures.
The Visualization (4) task is crucial, focusing on creating an output that maximizes users’ comprehension and engagement. This process revolves around transforming generated patterns, which inherently encapsulate the reified structure of input graphs, back into their original format — a multi-directed graph. In this graphical representation, relationships, generalizations, and associations manifest as edges, complete with corresponding cardinalities to provide a richer context.
The significance of this visual representation lies in its ability to empower users in several ways. For example, by translating the patterns from the FSM algorithm into multi-directed graphs, users can better grasp complex relationships and hierarchies within the data. This enables deeper insights and understanding. Figure 3 exemplifies the transformation power of visualization. On the left-hand side, you can observe a recurrent structure. On the right, you are presented with one of its occurrences, generated through the deepening task. This visual juxtaposition illustrates how .a can be transformed into a visually accessible and actionable representation, empowering users to extract valuable knowledge from their data with ease.
5 Implementation
We developed a command-line Python application for our approach, allowing users to interactively configure the process, manipulate the data, and assess the mining algorithm’s output. The scripts we created are available at https://github.com/unibz-core/CM-Mining.
For tasks involving graph importing, processing, and transformation, we used the NetworkX library444https://networkx.org/, a powerful Python package for creating, analyzing, and visualizing complex networks. NetworkX played a crucial role in each step of the entire workflow. We integrated the Grandiso library555https://pypi.org/project/grandiso/—a versatile Python library known for its efficient graph-matching and isomorphism-checking capabilities—in tasks involving graph matching, isomorphism detection, and motif search. For pattern mining, we relied on the gSpan Python implementation666https://pypi.org/project/gspan-mining/ of the frequent subgraph mining algorithm. This library enabled us to discover recurring subgraph patterns within our data. We achieved the visualization of discovered patterns using PlantUML777https://plantuml.com/, a flexible tool for generating UML diagrams. PlantUML seamlessly integrated into our framework, enhancing the interpretability of our results through clear and informative visual representations.
This suite of tools enabled a robust and efficient framework for structural pattern discovery and visualization. These tools collectively ensured the accuracy, scalability, and comprehensibility of our approach, enabling us to extract valuable knowledge from complex conceptual models.
5.1 Ontology-Driven Conceptual Modeling
To establish our scientific contribution’s proof of concept, we focused our implementation, in particular by adapting our importing step, on UML class diagrams, specifically on OntoUML models. These represent the output of a trending paradigm situated at the confluence of conceptual modeling and ontology engineering, namely Ontology-driven conceptual modeling (ODCM). ODCM frequently entails the utilization of fundamental ontologies to steer the formulation of conceptual models, modeling languages, and tools [34]. Within this framework, the OntoUML modeling language [35, 36], has risen to prominence as one of the predominant methodologies [34].
OntoUML extends the UML modeling language. Its meta-model is grounded on UFO (Unified Foundational Ontology) [37], a formal theory based on contributions from Formal Ontology in Philosophy, Philosophical Logic, Cognitive Psychology, and Linguistics. UFO is one of the most used foundational ontologies in conceptual modeling and OntoUML is among the most used languages in ontology-driven conceptual modeling [34]. An example of core ontological distinctions underlying OntoUML is represented by the key categories of object types (e.g., Kind, Subkind, Roles and RoleMixins), trope types (e.g., Relator, Mode) and Relations (FormalRelations, MaterialRelations and ParthoodRelations).888For an in-depth analysis and characterization of the ontological categories underlying OntoUML, the reader is referred to [35, 36].
One of the main goals of OntoUML is to support the conceptual modeling activities by making explicit the semantics behind the modelers’ design choices, thus enabling key features of the output conceptual models, such as understandability, interoperability, and reusability.
6 Experiments
In this section, we test our approach through three experiments, keeping as reference the requirements described in Section 3, and addressing the following research questions:
RQ1
Can the proposed approach identify previously recognized interesting patterns? This research question is aimed at testing whether the proposed solution can discover pre-identified interesting patterns (R1). This research question is also used to check the role of the customization steps in supporting the discovery process (R2) and the level of comprehensibility of the outcome (R3).
RQ2
What are the main parameters affecting the performance of the approach? Here, we want to identify the characteristics of the input data or parameters having a major influence on the performance (R4).
RQ3
Is the clustering step accurate in grouping patterns? This research question is mainly concerned with (R3). Here, we want to assess the practical utility of a key component in the presentation of the output.
The data about the experiments and detailed instructions for reproducibility are available at this GitHub link: unibz-core/CM-Mining_ExperimentData.
6.1 Experiment 1: Reliability Test
This experiment used an ad hoc set of models with some OntoUML patterns we already know as input and we checked how many expected patterns were found.
Data: A dataset comprising 10 models developed by our team. These models were created specifically for this experiment, taking inspiration from structures that are present in real models, with the goal of reproducing a controlled number of target patterns. The distribution of patterns per model was as from Table 1.
Setup: For validation, we used six common OntoUML patterns that served as “litmus test”.999A litmus test is “a critical indicator of future success or failure” A is a litmus test for B if A can be effectively used to measure some property of B. [38] These patterns were previously manually identified as useful for building OntoUML models by the designers of the language within multiple example models [7, 39]. We represent the selected patterns in Figure 4.
We executed the application seven times (one without adopting parameters to filter specific concepts and the other times with six different configurations, each to find one of the 6 selected target patterns) and we checked whether the proposed solution could discover the pre-identified interesting patterns (R1). Moreover, we tested the role of the customization steps in supporting the discovery process (R2) and the level of comprehensibility of the outcome (R3). First, we conducted a trial adopting no customization facility from the pipeline. As parameters, we selected 3 (the support of the less frequent patterns, i.e., Relator, Subkind and Phase) as minimum support, and 5 as minimum number of nodes (the size of the smaller pattern, i.e., Characterization),101010Where we have 2 nodes for the classes, 1 node as an association and 2 nodes for the cardinalities. so that all patterns can be discovered. For each other trial, we customized the pipeline to find the target pattern, e.g., by filtering some constructs and some other patterns we did not want to discover. For instance, for the Relator pattern, we selected Kind, Relator, and Role as target constructs and we filtered out information about the generalizations sets (the customization we adopted in terms of minimum support, nodes number and selected/removed constructs is reported in Table 2.).
Model | Relator | RoleMixin | Charact. | Category | Subkind | Phase |
01 | 0 | 0 | 1 | 1 | 1 | 1 |
02 | 0 | 0 | 2 | 1 | 0 | 1 |
03 | 0 | 0 | 1 | 1 | 0 | 2 |
04 | 1 | 0 | 1 | 0 | 0 | 0 |
05 | 1 | 0 | 1 | 1 | 0 | 0 |
06 | 0 | 1 | 0 | 2 | 0 | 0 |
07 | 0 | 1 | 1 | 0 | 1 | 0 |
08 | 1 | 0 | 1 | 1 | 0 | 0 |
09 | 0 | 1 | 0 | 2 | 0 | 0 |
10 | 0 | 1 | 1 | 0 | 1 | 0 |
Overall Freq. | 3 | 4 | 9 | 9 | 3 | 4 |
Model Freq. | 3 | 4 | 8 | 7 | 3 | 3 |
trial | freq. | nodes | filter | patterns |
Neutral | 3 | 5 | none | 4045 |
Relator | 3 | 12 | Select: ‘kind’, ‘role’, ‘relator’. Remove: ‘charact’ | 42 |
RoleMixin | 4 | 10 | Select: ‘kind’, ‘role’, ‘rmixin’, ‘genset’. Remove: ‘assoc’ | 2 |
Charact. | 8 | 4 | Select: ‘kind’, ‘mode’. Remove: ‘gen’ | 1 |
Category | 6 | 4 | Select: ‘kind’, ‘category’, ‘genset’. Remove: ‘assoc’ | 32 |
Subkind | 3 | 4 | Select: ‘kind’, ‘subkind’, ‘gen-set’. Remove: ‘assoc’ | 41 |
Phase | 3 | 4 | Select: ‘kind’, ‘phase’, ‘genset’. Remove: ‘assoc’ | 89 |
Results: The conducted trial identified the target patterns with 100% accuracy, demonstrating the pipeline’s proficiency in addressing R1. The approach identified all the (classes of) patterns that were explicitly identified a priori and correctly counted their occurrences.
The “Neutral” trial yielded an extensive list of output patterns, numbering in the thousands, augmenting the challenge of pinpointing the target pattern.
In subsequent trials, the application of the customization features significantly decreased the number of output patterns, thus showing the key role of the features we implemented in the discovery process (R2). For instance, with the Relator pattern, we selectively filtered extraneous stereotypes classes and relations out, resulting in a noteworthy reduction of the number of outputs (from 4045 to 42, see Table 2). We provide a visual representation and explain the Relator pattern in Figure 5. Particularly, the pipeline accurately and comprehensively identified all constructs originating from the input graphs.
Threats to validity: This experiment focused on a limited set of patterns that, although relatively intricate, represent only a fraction of a broader spectrum of patterns. Moreover, the synthetic dataset used in this study was fairly compact. However, the combination of the patterns we selected is suitable to assess whether our approach can discover structures with the level of expressiveness we require.
6.2 Experiment 2: Performance Test
Data: The input data comprised 94 models from a catalog of OntoUML models [14]. The catalogue results from a community effort, which collected models of different sizes, describing different domains, and developed for varying purposes in different contexts.
Setup: This validation was executed on a MacBook Pro (Retina, 13-inch, Early 2015) with CPU 2,7 GHz Intel Core i5 and was organized in two main trials: a baseline trial (0), where we selected as input the 47 conceptual models from the OntoUML dataset and trial (1) where we added 47 more models. The selected models have different sizes, in terms of relations and and classes. As an example, the graphs generated from these sources can have a small number of nodes and relations, namely 47 and 35 respectively (see lindeberg2022simple-ontorights.json, in the reference experiments repository), or higher, namely 2449 and 2282 (see indeberg2022full-ontorights.json). For each step, to increase the amount of data to be handled, we also tested the approach with six different customizations (we show these values in Table 3) concerning partition size parameters and the minimum frequency of the mining task.
Results: Table 3 summarizes the results of this experiment by providing the time taken by the graph partitioning, the mining step, the clustering, and the generation of a visual representation for each pattern. The execution time reveals that the processing time rises when:
.i
the number of models increases,
.ii
the number of nodes for the partitioned graph decreases, and
.iii
the minimum frequency threshold is decreased, thus allowing to find a larger number of patterns.
As we might expect, the function that takes the most time is the one dedicated to mining and pattern generation.
models | filtering (s) | nodes | freq. | mining (s) | patterns | clustering (s) | viz. (s) |
47 | 0,7327 | 5 | 20 | 1,375 | 1 | 0,011 | 0,106 |
3 | 20 | 1,591 | 15 | 0,043 | 1,648 | ||
5 | 15 | 3,792 | 22 | 0,073 | 2,257 | ||
3 | 15 | 3,950 | 65 | 0,422 | 6,669 | ||
5 | 10 | 47,579 | 246 | 0,982 | 16,387 | ||
3 | 10 | 48,448 | 372 | 1,360 | 20,387 | ||
94 | 1,238 | 5 | 20 | 130,458 | 37 | 0,103 | 3,781 |
3 | 20 | 184,239 | 72 | 0,503 | 7,360 | ||
5 | 15 | 224,782 | 141 | 0,734 | 9,410 | ||
3 | 15 | 229,138 | 193 | 0,812 | 11,500 | ||
5 | 10 | 324,781 | 586 | 1,990 | 25,675 | ||
3 | 10 | 359,318 | 662 | 3,156 | 31,057 |
Threats to validity: The primary challenge to the validity of this experiment stems from the absence of a comparative analysis involving different devices. The ultimate performance might be influenced by additional variables not present in the current setting. Nonetheless, the configuration we employed can be viewed as a stress test, given its resemblance to the common features found in widely utilized laptops and the absence of a high-performance CPU.
6.3 Experiment 3: Clustering Accuracy Test
Data: As input data, we used the same set of models used in experiment 1. In this experiment, we executed a mining task on the synthetic dataset adopted for experiment 1 and selected 33 out of the whole set of generated patterns, where each selected pattern can be easily traced back to one of the patterns presented in Figure 4 or one of its variations (e.g., Subkind patterns with a missing Subkind). We then manually clustered all the generated patterns by assigning a cluster label to them (e.g., we labeled the Subkind pattern and its variations as “cluster_0”). As the final segmentation, we produced six clusters, each one representing a pattern of Figure 4.
Setup: We applied our automated clustering step testing multiple similarity threshold values, e.g., 0.1, 0.15, 0.2, 0.25. We then compared the output of each test with the manually created dataset to calculate the accuracy of the automated clustering. Note that, in multiple cases, the number of clusters generated differed from those generated by hand (e.g., when the selected threshold was 0.1, the number of clusters was 2 against 6). For this reason, we calculated the accuracy as the percentage of correctly predicted cluster assignments out of the pairs of patterns being compared. More precisely, the function we adopted calculates the accuracy by comparing pairs of elements on two lists, namely: , the list of automatically clustered patterns, and , the list of manually clustered patterns. For each pair of patterns , the function checks whether pattern and in both lists and are inserted the same cluster. After checking all pairs , the function returns the accuracy as the ratio of correctly predicted pairs to the total number of pairs.
When the accuracy increased significantly, we decreased the distance of the threshold values, to understand what exactly was the point of best performance (for example, instead of going from 0.5 to 0.55, we tested 0.5, 0.51, 0.52, 0.53, 0.54, …, 0.64, etc.)
Results: We depict the outcomes in Figure 6. For the dataset we considered, the clustering component we employ in the pipeline achieves 0.93 accuracy when the adopted similarity threshold is 0.6 or 0.61. The increase in accuracy in this is due to the creation of the same number of clusters, such that the different patterns have been distinguished and the variations for each of them not deemed as different patterns. The approach incorrectly predicted only two patterns in the best case,, as shown in the confusion matrix in the figure.
Threats to validity: A primary concern regarding the validity of this experiment is that we conducted the test over a limited set of patterns by considering one clustering reference only. Users may vary in their way of classifying patterns, and it could be more challenging when the structures to be classified cannot be straightforwardly traced back to a reference pattern. Still, the test was enough to demonstrate how the clustering approach is reliable in distinguishing the structures that are generated and can be used to organize the data in a manner similar to what the user would adopt.
7 Utility Demonstration
To illustrate the applicability and utility of our proposed approach, we conducted a simulation of a use case. The primary objective of this simulation was to showcase the effectiveness of our approach in analyzing the practical usage of a particular modeling language: OntoUML. Our specific goal was to demonstrate how the discovery pipeline can be used to extract valuable insights that can enhance our understanding of the language, facilitate modifications, and drive improvements. Within this scope, we have used an extended version of the OntoUML dataset, comprising 143 models. To ensure a substantial number of patterns for analysis, we conducted five comprehensive tests, adjusting parameters as needed to avoid generating an overwhelming number of structures. In a second step, we went through the whole set of generated patterns and we engaged in internal discussions with two experts who contributed to the design of the language.111111Note that we have deliberately separated this section from the experiments section. The reason is that in this case we did not properly design an experiment, but simulated a use case with two language designers, who also participated in the requirements definition and design of the approach we are offering. That’s why we talked about “demonstration” and “internal” discussion. The interaction with the experts occurred in two stages. First, we presented the generated patterns and asked them to identify the patterns for which it was worth having a more in-depth discussion. Second, after defining a subset of 40 patterns, we went through a discussion of each of them, focusing primarily on .i the unexpectedness of the mined structure, and .ii the possibility of reusing the structure somehow. This simulation allowed us to address a new research question:
RQ4
To what extent can the approach be used to discover new interesting structures? This research question relates to R1, R2 and R3, and explores the approach’s capability to identify compelling structures—potentially unexpected ones—in real-world scenarios featuring a diverse set of models with varying levels of complexity. If identified, these structures can offer valuable insights for designers, potentially informing their future design strategies.
Next, we present four exemplary structures, taken from the subset of selected interesting patterns, which represent significant findings from our demonstration. Each example is accompanied by a brief discussion outlining the actions that could be undertaken based on our observations.
Example 1: In OntoUML, Modes are concepts representing particular types of properties with no structured values, which depend on their bearers [36]. According to the OntoUML specifications provided in [35] the constraints for modes are the following:
.i
every Mode must be (directly or indirectly) connected to an association end of at least one Characterization relation, and
.ii
the multiplicity of the characterized end (opposite to the Mode) must be exactly one.
Therefore, the situation in which a Mode is connected to two different bearers is not admitted.
Figure 7 shows a structure with three Modes, each one of them characterizing a distinct bearer, where two Modes are subtypes of a parent Mode.121212Nodes with dashed line denoted as class_0 or class_1, i.e., without associated stereotype, are nodes that in the discovered patterns present relationships about which we have information related to the source node, but not the target node, or vice versa. This is because, as mentioned above, the graphs given as input to the mining algorithm have all relationships reified. So, for example, in Figure 7 class_1 was created because in the pattern we had a mode node associated with an edge source to a relation characterization.
The information related to the occurrence of the pattern (.b) reveals that the reference berares are represented by a Role, a Category, and a Kind, respectively. This does not yet present an actual problem; however, it may involve the generation of scenarios that are not intended. In fact, the pattern suggests checking whether role and kind are subtypes of the same concept. In fact, if this is not the case, the child modes might end up associated with two different bearers (e.g., class_0 and class_1 may classify different instances). This specific situation may be avoided by forcing the bearers of the child modes (e.g., class_0 and class_2) to be subclasses of the bearer of the main parent mode. In conclusion, the recurring structure we found does not present an incorrect practice, but it may trigger an evolution of the language. Originally, in fact, the designers of OntoUML did not dwell on the possibility of constructing taxonomies of modes. However, through the results of this mining activity, emerges how the de facto modelers use these taxonomies (the importance of which was also discussed in [40]). This uncovered structure, then, opens the door to introducing new possible language constraints to avoid unwanted configurations.
Example 2: The structure provided by Figure 8 is significant for understanding how Events and their relationships are used in OntoUML.
Besides the structure we report here, many other variations having similar characteristics have been found. Consequently, a useful lesson we can learn from this is that the OntoUML module concerning events is less in control by modelers than that concerning, for example, Kinds or Roles. OntoUML has several relationships that can be used between Events. For example, see the common participation relationship or the HistoricalDependence relationships. The fact that the structure in Figure 8 and many similar ones do not have stereotypes for relations between Events can certainly be a sign that the users of the language do not properly understand these. In addition, by looking at the pattern occurrence information, one can see that some users have used relationship labels that invoke a specific stereotype (e.g., the case of ParticipationOf in the figure). In conclusion, this identified structure emphasizes the need for development guidelines for any related patterns for Event modeling.
Example 3: This example’s structure still relates to Event modeling. Here, what we represent in Figure 9 could be an occurrence of a syntactical mistake, depending on the cardinalities of the participation relation between event and RoleMixin. Here, the concept of RoleMixin [35] is the equivalent of Role, but instead of being played by instances of the same Kind (e.g., student-person), it is played by instances of different Kinds. Examples of RoleMixins include customers and insured items, the former being played by both people and companies, whilst the latter being played by cars, houses, and paintings [41].
That being considered, the OntoUML RoleMixin construct requires that there are at least two identity providers involved and then at least two different Role instances. Thus, in what sense does a RoleMixin participate in an Event? Looking at the occurrence of the pattern, the concept of RoleMixin is modeled as “client” and in this sense, we may have different instances of this class that are, for instance, persons or organizations. Considering this structure, to avoid possible mistakes, the cardinalities of the relationship between the RoleMixin concept and the Event concept must be made explicit. The recurrent structure we found indicates that the modeler left this information implicit. Consequently, the possible insight that can be collected by the language designer through the assessment of this structure is that the modeler should be guided in making explicit the connection between the identity providers and the given Event, possibly by providing the right cardinalities of the associations in question.
8 Discussion
RQ1: Can the proposed approach identify previously recognized interesting patterns? The answer to this question is affirmative. Experiment 1 demonstrated that already-known structures can be discovered accounting for all the constructs of which they are composed and without missing any occurrence. For the patterns we selected, we discovered all classes of patterns and their occurrences with 100% accuracy. The patterns we have chosen were also useful to demonstrate how the approach can account for the core constructs in the input language. The experiment also highlighted the importance of customization options is important in reducing the number of results, thus facilitating the identification process. The importance is enlarged when the methodology is applied to extensive real-world datasets, such as the one represented by the OntoUML catalog.
RQ2: Which are the main parameters affecting the performance of the proposed approach? According to Table 3, the increase in pattern size and minimum support value improves the average performance, but decreases the number of patterns that can be found. Unlike when the goal is to search for more frequent structures, we can conclude from this that when it is necessary to find more information or possibly unexpected information, it is necessary to find the right trade-off between effectiveness and performance. On a dataset comprising OntoUML models, if we search too small and too infrequent patterns, the algorithm can produce even more than 10,000 outputs, often very similar to each other. Here, too, the role of customization steps can be essential, especially in reducing the information in the graphs to be sent as input to the mining algorithm and in eliminating redundant or unwanted outputs, thus also reducing the time of visualization and assessment.
RQ3: Is the clustering step accurate in grouping patterns? Thanks to Experiment 3, we observe how our proposed clustering component provides a grouping criterion for patterns akin to manual user-performed ones. Our current configuration enables the differentiation of various structures by considering both the node count and the types of constructs involved. This capability positions it as a valuable asset during the results assessment phase. However, a point of discussion centers on the heterogeneity of the threshold values needed to achieve a satisfactory grouping, which, as we have found in other tests, can vary from case to case. We cannot assert a universally optimal threshold, such as adhering to a specific value like 5, for attaining the best grouping. Our trials have shown that optimal thresholds may range around 0.6 or 0.7 in some instances. One consistent observation, though, is that the optimal range for classification consistently falls within the mid-range of the scale. That is, thresholds between 0.1 and 0.3 or between 0.8 and 1.0 do not consistently achieve the most effective clustering. Considering the iterative nature of the discovery process, where users often experiment and fine-tune parameters to get the desired output in terms of size and type of structures, adjusting threshold values to achieve the desired clustering is also part of the iterative discovery process.
RQ4: To what extent is the approach useful for discovering new interesting structures? From the tests we did for the utility demonstration, we observed that to find unexpected modeling practices, the importance value of frequency (or minimum support) decreases. In fact, if parameters are set to find frequent structures, it is more likely that we will find already known structures, especially with a language like OntoUML, where modelers operate following predefined guidelines and patterns. This mainly has a negative effect on the overall performance of the approach. In fact, as we saw in Experiment 2, the more we decrease the frequency threshold, the longer the mining process takes. Moreover, the approach has the potential to generate patterns of a very large order of magnitude. However, this did not prevent us from collecting useful and interesting information, such as that discussed in the previous section, which can trigger a range of analyzes and possibly interventions by the language designers that they had not thought of before. Currently, our emphasis is not on exploring the interesting types and unexpected structures that can be found, as it is not the main purpose of the paper. Our understanding of classifying these outputs is growing as we consider the analyzes and operations they might trigger.
Structures triggering the definition of new patterns or anti-pattern. For instance, structures like the one represented in Figure 7 could be a good start for understanding and, possibly, devising new modeling strategies to be suggested to modelers. In the specific case of the example given, the structure of modes is a concrete case for understanding how to enable the designer to avoid generating unintended instantiations. For example, designers can concretely achieve this by exploiting the found structure and offering an ideal way of modeling modes taxonomies, or by introducing new constraints that force the creation of generalization links between the concepts characterized by the sub-modes and the concept characterized by the parent-mode.
Structures triggering a clarification of constructs and their usages. A typical example in this regard is that represented by Figure 8. Here, the pervasive absence of relation stereotypes composing structures involving events is a clear sign of how the modeling guidelines related to the area of the language devoted to these constructs (in this case, the area dependent on UFO-B [42]) can be further elaborated.
Structures that highlight possible adoptions of bad practices. In this sense, the approach is useful in unearthing new types of errors or understanding whether practices deemed to be avoided are actually adopted, and to what extent. The example provided by Figure 9 is a typical case and demonstrates the utility of the approach since it can be a strong ally in understanding how to possibly guide modelers to not omit key information or extend the language with new constraints.
9 Related Work
There is extensive literature [43, 44, 45, 29] on pattern discovery and its applications in a variety of domains, including software code [46], databases [47], educational processes [48], business processes [49], etc. However, the use of pattern discovery techniques is notably restricted in conceptual models. In this focused area of research, the closest work to what we propose is that of Skouradaki et al. [8], who designed a pattern mining algorithm for BPMN. Still, the goal of our contribution is not to provide a new mining algorithm. Our emphasis predominantly lies in the combination of a well-established FSM technique with graph manipulation techniques. Furthermore, a considerable amount of effort from our side concerns the definition of an interactive process where users can participate in the discovery activities, thus affecting the reliability of the final output. Last, we designed the approach with the scope of covering different conceptual modeling languages by keeping all the functions of the approach as language-independent.
Ławrynowicz et al. [12] seek to discover domain patterns related to specific areas of information and independent of the modeling language constructs that recur across OWL ontologies by applying a tree-mining technique. They divide their contribution into two main steps, which partially resemble aspects of our strategy, namely: a transformation step, where ontology axioms are transformed into tree structures; and an association analysis step, where co-occurring axioms are extracted to discover ontology patterns. This research is applied to a set of ontologies from the BioPortal repository and is very similar to ours in spirit. However, our solution presents key differences. First, for the mining step, we adopted the frequent subgraph mining algorithm, thus involving a completely different input preparation step. Second, we devised our approach with the main goal of discovering structural modeling patterns, namely patterns defined simply by the combination of constructs of a modeling language. In Ławrynowicz et al.’s work, the discovered patterns concern primarily domain-specific information that may recur within or across ontologies (e.g., what are the recurrent properties of the class “person”) [12]. Again, the interaction capabilities we proposed are out of their scope.
In the same direction, Lee et al. [10] seek to discover domain patterns across and within ontologies. However, to address this challenge, two different steps are adopted: a step where sub-graphs are extracted through candidate generation and chunking processes; a step where frequent sub-graphs mining [50] is adopted. This work also focuses on domain-specific patterns and one of its priorities is to allow the processing of large-scale knowledge graphs. Furthermore, the paper lacks instructions on how to handle an interactive discovery process.
Two other approaches that are worth mentioning are the ones presented by Ardimento et al. [51] and the one that was introduced by Fumagalli et al. [52]. Both refer directly to UML class diagrams as the reference modeling language to be mined. However, the former is more aimed at mining modeling events that occur on Visual Paradigm,131313https://www.visual-paradigm.com/download/community.jsp i.e., the reference editing tool, and not properly frequent subgraph structures. The second, on the other hand, is much closer to our approach and is primarily focused on using frequent item set mining, instead of frequent subgraph mining, and also does not introduce the set of facilities we have offered here to support the discovery process.
To summarize, our approach distinguishes itself from related work through two key novel contributions. First, we empower users in their discovery process by offering a comprehensive set of interactive steps, providing guidance and assistance throughout the discovery journey. Second, we employ frequent subgraph mining, a robust established mining technique, for mining subgraphs in highly complex UML class diagrams represented by the models hosted in the OntoUML dataset. Our approach accounts for all vital elements of these diagrams, including cardinalities, generalization sets, multiple edge labels and class labels, and multi-directionality, thus ensuring the possibility of finding patterns containing all this information.
10 Final Considerations
This paper presents a practical interactive approach, whose implementation is available at https://github.com/unibz-core/CM-Mining, for automating the empirical discovery of modeling patterns in conceptual models by combining state-of-the-art graph manipulation techniques and frequent subgraph mining. Structural patterns are reusable representations of bad or good practices. They are typically harvested by manual labor-intensive processes, which usually take a long time to converge into pattern catalogs. Our automated process for pattern discovery facilitates the construction of pattern catalogs tailored for modeling languages. Additionally, we create a mechanism for helping language designers to create higher-granularity primitives in that language, i.e., modeling patterns that can become part of the grammar and tools of that language [7].
Based on the encouraging results presented here, we created a list of tasks for immediate future work. First, we are going to further assess the approach to finding unexpected patterns. In order to address this, we will involve representative users and run an evaluation through usability testing. Second, we are going to test the approach with models encoded in different conceptual modeling languages, such as BPMN or ArchiMate (R5). In the current work, we used a specific subset of (Onto)UML as a target use case scenario, leveraging a large dataset of high-quality data. However, we have designed our solution to rely on graph-structured data, enabling the encoding of any conceptual model construct. Moreover, the whole approach, which can be adapted to different input formats, is completely language-independent except for the first importing step.